In a certain liberal arts college with about 10,000 students, 47% are males. If two students from this college are selected at random, what is the probability that they are of the same gender? Your answer should be rounded to 4 decimal places.

According to the information that comes with a certain prescription drug, when taking this drug, there is a 22% chance of experiencing nausea (N) and a 48% chance of experiencing decreased sexual drive (D). The information also states that there is a 14% chance of experiencing both side effects.

What is the probability of experiencing only nausea?  

Your answer should be rounded to two decimal places.
Answer #1

Total number of students, n= 10000

1. P( males) = 0.47

P( Females) = 1-0.47 = 0.53

P( Both are of same Grender) = P( Both are male) + P( Both are Female)

= ( 0.47)2 + ( 0.53)2

= 0.5018

2. P( Nausea) = P( N) =0.22

P( Decreased sexual derive) = P(D) =0.48

P( N n D) = 0.14

P( Only Nausea) = P( only N) = P( N) - P(N n D) = 0.22 - 0.14 = 0.08